A grinding wheel is in the form of a uniform solid disk of radius 6.93 cm and mass 2.02 kg. It starts from rest and accelerates uniformly under the action of the constant torque of 0.597 N · m that the motor exerts on the wheel. How long does the wheel take to reach its final operating speed of 1300 rev/min? Through how many revolutions does it turn while accelerating?
Answer #1

The moment of inertia of the grinding wheel about its axis of rotation is given by, mr 2 (2.02 kg) (6.93x10 m) 4.85x10-3 kg m
Answer #2

T = Iα = mr2α/2

=> α = 2T/mr2 = 2 * 0.597 / (2.02 * 0.06932) = 123.08 rad/s2

Time taken, t = ω/α = mr2ω/2T = (1300 * 2π) / 123.08 = 66.4 s

Revolutions turned, θ = (1/2π)(αt2/2) = αt2/4π = 123.08 * 66.42 / 4π = 43183 revolutions